OpenStudy (exoexoexo):

the cost C for maintenance on a new hybrid car increases each year by 10%. if frank paid $400 this year for maintenance, then the cost x years from now can be given by the function C(X)= 400(n)^x what is the value of n?

OpenStudy (exoexoexo):

\[C(X)= 400(n)^{x}\]

jimthompson5910 (jim_thompson5910):

x = number of years C(x) = the cost for some year x n = growth factor the growth in costs is 10% so the growth factor is 1+10% = 1+0.1 = 1.1 making n = 1.1 and the C(x) equation to be \[\Large C(x) = 400(1.1)^x\]

jimthompson5910 (jim_thompson5910):

By the way, 10% is really high for cost growth. Either Frank should find another car or find someone else to fix his car (that has cost growth less than the rate of inflation).